### v1

@@ 5 Comparison with previous work and improved scaling relations
In this section, we show that the new measurements are consistent with the previous ones (Figure 9), with no average offset. We compared the BIC-weighted velocity dispersion of 37 galaxies with the corresponding measurements in Sonnenfeld et al. (2013b) for which the same spectra were used. For other cases, different spectra were used, or no measurements were reported. We found that for the 37 galaxies, the estimated $<y_{i}/x_{i}>=0.98\pm 0.02$ where $y_{i}$ and $x_{i}$ are new and old measurements for the $i$th galaxy. Thus, the current velocity dispersions are consistent with the previously measured values on average. The reduction in systematic error with respect with the previous analysis is due mostly to three factors. First, stellar template libraries have been improved. Whereas Sonnenfeld et al. (2013b) used 7 stars from the Indo-US library, we use hundreds of stars from the cleaned Indo-US and MILES libraries. Second, we do not use the data taken with instrumental setups that have insufficient spectral resolution to meet our target accuracy. Third, we have mitigated the effects of continuum and flux calibration uncertainties by carefully modeling them with multiplicative and additive polynomials.
@@
The distribution of velocity dispersions in the SL2S sample does not vary significantly with redshift (as we show in Figure 7), and peaks at significantly higher redshift than SLACS ($z=0.1-0.4$).
@@
Figure 9: Comparison of velocity dispersions for 37 galaxies as measured in this paper (y-axis) and by Sonnenfeld et al. (2013a, x-axis) based on the same spectra. The only thing that changed is the methodology, chiefly improvements in the stellar template libraries.
@@
We then use the new SL2S measurements to revisit the scaling relations determined by Auger et al. (2010) for the SLACS sample, taking advantage of the new spectroscopy determinations by Knabel et al. (2024) and those presented here. Note that to enable direct comparison to Auger et al. (2010), the stellar velocity dispersion has been corrected to a fiducial aperture of half the effective radius using the empirical recipes derived by Knabel et al. (2024). We refer to these aperture corrected velocity dispersions as $\sigma_{1/2}$. We adopt total uncertainties as the sum in quadrature of the random and systematic errors.
Our goal is to investigate whether there are any differences between the SLACS, SL2S, and TDCOSMO samples (Millon et al. 2020; Birrer et al. 2020, and TDOCOSMO collaboration 2025, in preparation) that may potentially be an indication of structural evolution or differences in terms of these properties.
@@
We limit ourselves to the SLACS and SL2S galaxies with high SNR spectroscopy from Knabel et al. (2024, 2025) and this work, to minimize the potential impact of systematic errors. We also limit ourselves to the scaling relations that involve lensing quantities, as the others can be investigated using much larger samples of non-lens galaxies. Effective radii and Einstein radii are taken from Auger et al. (2010) for the SLACS sample, from Sheu et al. (2025; submitted) for the SL2S sample, and from TDCOSMO collaboration (2025, in prep) for the TDCOSMO sample. The effective radii are obtained by fitting Sérsic (1963) profiles to the surface brightness of the deflector, jointly with the lensed images444de Vaucouleurs (1948) in the case of SLACS-X. Noting that the formal errors are usually tiny on these quantities, we adopt 10% as our systematic error on the effective radius (e.g., Cappellari et al. 2013a, sec. 4.1.1) and 2% on the Einstein radius (e.g., Bolton et al. 2008a). We neglect possible contributions to the lensing quantities by external convergence, since it is at a few percent level at most (Birrer et al. 2020) and would not affect our conclusions in a significant way.
@@
In addition to the scaling relations described by Auger et al. (2010), we consider a new correlation that involve the mass density profile slope $\gamma_{\rm lens}$ determined from lens models (Tan et al. 2024, and Sheu et al. 2025, submitted). This quantity was not available to Auger et al. (2010). We note that Auger et al. (2010) used a slope $\gamma$ derived by a joint lensing and dynamical analysis, which is the same only if the mass density profiles are exactly power laws (Sonnenfeld et al. 2013b).
All the quantities presented in this analysis are summarized in Table 3.
@@
5.1 Correlation between lensing mass profile slope and ratio between lensing and stellar velocity dispersion
@@
The first correlation that we examine is one between lensing slope and the ratio between the stellar velocity dispersion $\sigma$ and lensing velocity dispersion $\sigma_{\rm SIE}$, $f_{\rm SIE}=\sigma/\sigma_{\rm SIE}$. The lensing velocity dispersion is related to the Einstein radius $\theta_{\rm E}$ and the angular diameter distance to the source (Ds) and between the deflector and the source (Dds) by the equation (Treu 2010):
@@
The subscript SIE is used as a reminder that the Einstein radius is usually derived by fitting a singular isothermal ellipsoid to the lensing features. It is worth noticing however that the Einstein radius is very robust to the choice of lens model and only varies by 1-2% if a power law model is adopted instead (Bolton et al. 2008b).
Treu et al. (2006) introduced $f_{\rm SIE}$ as a proxy for the slope of the mass density profile, finding that it should be close to unity, as expected for an isothermal mass density profile. This can be understood qualitatively in terms of spherical profiles. Given a power law total mass density profile with three-dimensional form $\rho_{\rm tot}\propto r^{-\gamma}$, the mass enclosed within a two-dimensional circular aperture with radius $R$ is given by
@@
where RE is the Einstein radius in kpc and ME is the mass within the Einstein radius. Following Schneider et al. (2006), it is easy to show that
@@
where Reff is the effective radius and $c_{\rm e2}$ is a dimensionless structural parameter than depends on the three dimensional shape of the mass distribution, orbital structure, and spatial distribution of luminous tracers (Nipoti et al. 2008). Bolton et al. (2008b) finds $\langle\log c_{\rm e2}\rangle=0.53\pm 0.06$. In the following we will assume that ce2 is a constant, which is a good approximation. Residuals trends from the predicted relations would imply that ce2 is not exactly a constant, but that it depends, e.g., on galaxy mass, stellar velocity dispersion, and mass density profile slope.
By combining equations 9 to 11 one finds that
@@
In Figure 10 we compare the measured fSIE with the values predicted by Equation 12, using ce2 measured by Nipoti et al. (2008). Even though the data are noisy, we note that the SLACS-KWCI and TDCOSMO samples are very close to the predicted values. There is more scatter for the SL2S, which is in part expected owing to the fact that $R_{\rm eff}\sim R_{\rm E}/2$, and thus the factor to be elevated to the power of $2-\gamma_{\rm lens}$ is on average 4, introducing extrapolation errors. Furthermore, ce2 is not a constant and is expected to depend significantly on $\gamma$. Thus using the average value will be accurate for systems with $\gamma\sim 2$, similar to the SLACS sample and less so if $\gamma$ departs from 2.
Figure 10: Comparison between the measured $f_{\rm SIE}$ and that derived from Equation 12.
5.2 The lensing mass fundamental planes
Bolton et al. (2007) introduced the lensing mass fundamental plane (MFP; see also Bolton et al. 2008b; Auger et al. 2010) by analogy with the standard fundamental plane, i.e. replacing surface brightness with surface mass density:
@@ 5.2 The lensing mass fundamental planes
This MFP could provide some useful insight. For a power law, the surface mass density within a projected radius is given by
@@
where $\Sigma_{c}$ is the lensing critical density
@@
where $\Sigma_{1/2}$ is the average surface mass density within half the effective radius. Taking the logarithm of both sides and sorting the variables yields
@@
The coefficients are found to be close to $a=2$ and $b=-1$ respectively, indicating that the lensing mass fundamental plane is much better described by this simple interpretation than the classic fundamental plane, which appears to be ”tilted” with respect to the virial prediction, requiring variations in M/L ratio with $\sigma$, generally attributed to changes in dark matter content, stellar population age, metallicity, and initial mass function (Treu et al. 2006, 2010; Auger et al. 2010). The intercept of course depends on the units. Consistent results, with coefficients in agreement with virial predictions, are obtained when the mass fundamental plane is measured using stellar dynamics fitted to integral-field stellar kinematics (e.g., Cappellari et al. 2006, 2013a; Zhu et al. 2024).
@@
Previous studies (Bolton et al. 2007, 2008a; Auger et al. 2010) computed the lensing mass fundamental plane assuming isothermal mass profiles ($\gamma=2$). In Figure 11 we derive the MFP under the same assumption, using the SL2S sample, SLACS-KCWI sample, and the SLACS sample reported by Auger et al. (2010) based on SDSS velocity dispersions (SLACS-X). To fit the plane, we use Python package LtsFit (Cappellari et al. 2013b), a robust least squares regression with uncertainties in all dimensions and intrinsic scatter. We do not clip outliers but include all datapoints in the fit. Effective radii are fit in kpc, velocity dispersions in units of 250 $\mathrm{km\ s^{-1}}$, and surface mass densities in $\mathrm{10^{10}\ M_{\odot}\ kpc^{-2}}$. The samples follow the same MFP with coefficients consistent with those found by Auger et al. (2010). The slopes are close but not identical to those in Equation 17, indicating that the assumption of a constant virial coefficient is not exactly correct, but it is likely to have some dependence on mass or velocity dispersion, as elliptical galaxies are not completely scale invariant. We overplot the TDCOSMO sample for illustration, showing them to lie along the plane across the sample space occupied by the SL2S and SLACS samples. The model indicates no intrinsic scatter to the combined sample.
@@
Figure 11: Isothermal Mass Plane (MP) $\mathrm{\log R_{eff}=a\log\sigma_{1/2}+b\log\Sigma_{1/2,iso}+c}$. Plane is fitted to the combined SL2S, SLACS/KCWI, and SLACS-X samples. TDCOSMO sample is overplotted. Here $\mathrm{\Sigma_{1/2,iso}=M_{R_{eff}/2,iso}/\pi(R_{eff}/2)^{2}}$ is the average lensing surface mass density within half the effective radius assuming the mass density profile is isothermal (i.e., $\gamma_{lens}=2$). $\sigma_{1/2}$ is given in units of 250 $\mathrm{km\ s^{-1}}$, $\mathrm{M_{1/2}}$ in units of $10^{10}\ \mathrm{M_{\odot}}$, and $\mathrm{R_{eff}}$ in kpc. As a test, we fitted all combinations of subsamples, and the coefficients are the same within the errors. The dotted line is the one-to-one line.
@@
We are also in a position to study for the first time the MFP for a generic power law profile, using the measurements of $\gamma=\gamma_{\rm lens}$ that are now available. The uncertainty on $\Sigma_{1/2}$ is given by
@@
$\delta\log\Sigma_{1/2}=\sqrt{(1-\gamma_{\rm lens})^{2}\left(\frac{\delta R_{\rm eff}}{R_{\rm eff}\ln 10}\right)^{2}+\delta\gamma^{2}\left(\log\frac{R_{\rm eff}}{2R_{E}}\right)^{2}}$ (18)
The resulting MFP is shown in Figure 12. As expected, the errors are larger than in the isothermal version, especially for the SL2S sample. For this sample, $\delta\gamma_{\rm lens}$ is generally larger than for TDCOSMO and SLACS due to the lower SNR of the imaging data. Furthermore, the Einstein radius is typically several times the effective radius, requiring more extrapolation than for the other samples. This plane is consistent with no scatter within its error bars as well, but in Figure 12 we report the best fit, which includes 1% intrinsic scatter.
@@
Both planes are well defined, and there are no significant differences between the sub-samples. We quantify any residual differences between the sub-samples in the following way. Since the sub-samples, except SLACS-X, are too small to robustly determine the slopes, we keep the slopes $a$ and $b$ fixed to the values obtained for the combined sample and then we fit for the intercept $c$ of each subsample. In the isothermal case, the intercept is $0.451\pm 0.019$, $0.463\pm 0.027$, $0.456\pm 0.027$, and $0.445\pm 0.015$, respectively for the combined, SL2S, SLACS-KCWI, and SLACS-X samples. They are consistent within the errors. The error on the intercept is largest for the SL2S sample and corresponds to $\pm$5.8in Reff and ±$\pm 4.2\%$ in $\sigma_{1/2}$. For the non-isothermal case, the intercept is $0.560\pm 0.020$, $0.579\pm 0.029$, $0.561\pm 0.021$, and $0.548\pm 0.022$ for the samples in the same order as above. Again, they are consistent within the errors. The error is largest for the SL2S sample and corresponds to $\pm$6.9in Reff and ±$\pm 5.3\%$ in $\sigma_{1/2}$.
@@
Figure 12: Mass plane (MP) $\mathrm{\log R_{eff}=a\log\sigma_{1/2}+b\log\Sigma_{1/2}+c}$. The plane is fitted to the combined SL2S, SLACS/KCWI, and SLACS-X samples. TDCOSMO sample is overplotted. Here $\mathrm{\Sigma_{1/2}=M_{R_{eff}/2}/\pi(R_{eff}/2)^{2}}$ is the average lensing surface mass density within half the effective radius. We have updated the SLACS-X lens models to those from Tan et al. (2024) with free power law slopes to compare with SL2S and SLACS/KCWI lens models. $\sigma_{1/2}$ is given in units of 250 $\mathrm{km\ s^{-1}}$, $\mathrm{M_{1/2}}$ in units of $10^{10}\ \mathrm{M_{\odot}}$, and $\mathrm{R_{eff}}$ in kpc. The dotted line is the one-to-one line.
@@
In conclusion, the MFP analysis shows no evidence of structural differences between the SL2S, SLACS, and TDCOSMO samples. This does not mean that there is no evolution in the characteristics of the galaxies. For example, in Figures 11 and 12 the SL2S galaxies tend to have smaller effective radii, generally consistent with the finding that early type galaxies have smaller sizes at high redshift that in the local universe for a given stellar mass (e.g., van Dokkum et al. 2010). Sonnenfeld et al. (2015) found that the fraction of dark matter within the inner 5 kpc increases with redshift, and Sheu et al. (2025, submitted) also found mild evidence for evolution. However, this work shows that galaxies evolve within the lensing mass plane. An example of a mechanism that moves galaxies along the lensing mass plane is the ’envelope accretion’ recently considered by Nipoti (2025) in addition to the standard major and minor mergers. Envelope accretion describes the case in which diffuse satellites are completely disrupted in the galaxy outskirts. As show in Figure 13, tantalizingly, the model proposed by Nipoti (2025) predicts that envelope accretion makes galaxies evolve along the lensing mass plane, almost exactly. Similar results are obtained for the generalized mass plane with free mass density slopes. A full comparison with galaxy evolution models is beyond the scope of this paper and is left for future work.
Figure 13: Evolution of a reference galaxy in the lens mass fundamental plane according to the ”envelope accretion” model by Nipoti (2025). The initial (red symbol) and final (blue symbol) positions of the galaxy in the plane are calculated using the model described in section 6.2 of Nipoti (2025), in which the galaxy grows by a factor of two in stellar mass.

### v2

@@ 5 Comparison with previous work and improved scaling relations
In this section, we show that the new measurements are consistent with previous results (Fig. 9), with no average offset. We compared the BIC-weighted velocity dispersion of 37 galaxies with the corresponding measurements by Sonnenfeld13b, for which the same spectra were used. For other cases, different spectra were used, or no measurements were reported. We found that for the 37 galaxies, the estimated $y_{i}/x_{i}>=0.98\pm 0.02$ where $y_{i}$ and $x_{i}$ are new and old measurements for the $i$th galaxy. Thus, the current velocity dispersions are consistent with the previously measured values on average. The reduction in systematic error with respect to the previous analysis is mostly due to three factors. First, stellar template libraries have been improved. Whereas Sonnenfeld13b used 7 stars from the Indo-US library, we used hundreds of stars from the cleaned Indo-US and MILES libraries. Second, we did not use the data taken with instrumental setups that have insufficient spectral resolution to meet our target accuracy. Third, we mitigated the effects of continuum and flux calibration uncertainties by modeling them with multiplicative and additive polynomials.
@@
The distribution of the velocity dispersions in the SL2S sample does not vary significantly with redshift (as we showed in Fig. 7), and it peaks at a significantly higher redshift than SLACS (z=0.10.4z=0.1-0.4).
@@
Figure 9: Comparison of the velocity dispersions for 37 galaxies as measured in this paper (y-axis) and by Sonnenfeld_2013a based on the same spectra. The galaxies are color-coded by the S/N measured in this work. The method alone changed, i.e., chiefly improvements in the stellar template libraries.
@@
We then used the new SL2S measurements to revisit the scaling relations determined by Auger_2010 for the SLACS sample, taking advantage of the new spectroscopy determinations by Knabel24 and those presented here. To enable a direct comparison with Auger10, the stellar velocity dispersion was corrected to a fiducial aperture of half the effective radius using the empirical recipes derived by Knabel24. We refer to these aperture-corrected velocity dispersions as $\sigma_{1/2}$. We adopted the total uncertainties as the sum in quadrature of the random and systematic errors.
@@
Our goal was to investigate whether there are any differences between the SLACS, SL2S, and TDCOSMO samples (Millon_2020_TDCOSMO_I; birrer20; Milestone25) that might be an indication of structural evolution or differences in terms of these properties. We focused on the SLACS and SL2S galaxies with high S/N spectroscopy from Knabel24 and this work to minimize the potential impact of systematic errors. We also focused on the scaling relations that involve lensing quantities because the others might be investigated using much larger samples of nonlens galaxies. The effective radii and Einstein radii were taken from Auger_2010 for the SLACS sample, from William_2025_dinos_2 for the SL2S sample, and from Milestone25 for the TDCOSMO sample. The effective radii were obtained by fitting Sersic_profile profiles to the surface brightness of the deflector, jointly with the lensed images555deVaucouleurs48 in the case of SLACS-X. Noting that the formal errors are usually tiny on these quantities, we adopted 10% as our systematic error on the effective radius (e.g., ATLAS_XV_Michele_2013, sec. 4.1.1) and 2% on the Einstein radius (e.g., Bolton08). We neglected possible contributions to the lensing quantities by external convergence because it is at a few percent level at most (birrer20) and would not affect our conclusions in a significant way.
@@
In addition to the scaling relations described by Auger_2010, we considered a new correlation that involves the mass-density profile slope $\gamma_{\rm lens}$ determined from lens models (T24; William_2025_dinos_2). This quantity was not available to Auger_2010. We note that Auger_2010 used a slope $\gamma$ derived by a joint lensing and dynamical analysis, which is the same only when the mass-density profiles are exactly power laws (Sonnenfeld13b). All the quantities presented in this analysis are summarized in Table 3.
@@
5.1 Correlation between the lensing-mass profile slope and ratio of the lensing and stellar velocity dispersion
@@
The first correlation that we examined is the correlation between the lensing slope and the ratio of the stellar velocity dispersion $\sigma$ and the lensing velocity dispersion $\sigma_{\rm SIE}$, $f_{\rm SIE}=\sigma/\sigma_{\rm SIE}$. The lensing velocity dispersion is related to the Einstein radius $\theta_{\rm E}$ and the angular diameter distance to the source (Ds) and between the deflector and the source (Dds) by the equation (Treu10b)
@@
The subscript SIE is used as a reminder that the Einstein radius is usually derived by fitting a singular isothermal ellipsoid to the lensing features. It is worth noting, however, that the Einstein radius is very robust to the choice of lens model and only varies by 1-2% when a power-law model is adopted instead (Bolton08b).
Treu06 introduced $f_{\rm SIE}$ as a proxy for the slope of the mass density profile, finding that it should be close to unity, as expected for an isothermal mass density profile. This can be understood qualitatively in terms of spherical profiles. Given a power-law total mass-density profile with the three-dimensional form $\rho_{\rm tot}\propto r^{-\gamma}$, the mass enclosed within a two-dimensional circular aperture with radius $R$ is given by
@@
where RE is the Einstein radius in kpc, and ME is the mass within the Einstein radius. Following Schneider06, it is easy to show that
@@
where Reff is the effective radius, and $c_{\rm e2}$ is a dimensionless structural parameter that depends on the three-dimensional shape of the mass distribution, orbital structure, and spatial distribution of luminous tracers (NTB08). Bolton08b finds $\langle\log c_{\rm e2}\rangle=0.53\pm 0.06$. In the following, we assume that ce2 is a constant, which is a good approximation. Residual trends from the predicted relations would imply that ce2 is not exactly a constant, but that it depends on galaxy mass, stellar velocity dispersion, and the mass-density profile slope, for instance.
By combining Eqs. ( 9) to ( 11), we find that
@@
In Fig. 10 we compare the measured fSIE with the values predicted by Eq. ( 12), using ce2 measured by NTB08. Even though the data are noisy, we note that the SLACS-KCWI and TDCOSMO samples are very close to the predicted values. There is more scatter for the SL2S, which is in part expected because $R_{\rm eff}\sim R_{\rm E}/2$, and thus, the factor to be elevated to the power of $2-\gamma_{\rm lens}$ is 4 on average, introducing extrapolation errors. Furthermore, ce2 is not a constant and is expected to depend significantly on $\gamma$. Thus, using the average value is accurate for systems with $\gamma\sim 2$, similar to the SLACS sample and less so when $\gamma$ departs from 2.
Figure 10: Comparison between the measured $f_{\rm SIE}$ and that derived from Eq. ( 12).
5.2 The fundamental planes of the lensing mass
Bolton07 introduced the lensing-mass fundamental plane (MFP; see also Bolton08b; Auger_2010) by analogy with the standard fundamental plane, that is, they replaced the surface brightness with the surface-mass density,
@@ 5.2 The fundamental planes of the lensing mass
This MFP could provide some useful insight. For a power law, the surface-mass density within a projected radius is given by
@@
where $\Sigma_{c}$ is the critical surface mass density
@@
where $\Sigma_{1/2}$ is the average surface-mass density within half the effective radius. Taking the logarithm of both sides and sorting the variables yields
@@
The coefficients are found to be close to $a=2$ and b=1b=-1, indicating that the virial coefficient is close to a constant, and thus, the lensing-mass fundamental plane is much better described by this simple interpretation than by the classic fundamental plane, which appears to be tilted with respect to the virial prediction. This requires variations in the M/L ratio with $\sigma$, which are generally attributed to changes in the dark matter content, stellar population age, metallicity, and initial mass function (CLR96; TBB04; Dutton13; Bernardi20; Deugenio21; Treu06; Treu10; Auger10). The intercept depends on the units. Consistent results, with coefficients in agreement with virial predictions, are obtained when the mass plane (MP) is measured using mass from stellar dynamics along with integral-field stellar kinematics (e.g., Cappellari2006; ATLAS_XV_Michele_2013; Zhu2024).
@@
Previous studies (Bolton07; Bolton08; Auger_2010) computed the lensing-mass fundamental plane assuming isothermal mass profiles ($\gamma=2$). In Fig. 11 we derive the MFP under the same assumption, using the SL2S sample, SLACS-KCWI sample, and the SLACS sample reported by Auger_2010 based on SDSS velocity dispersions (SLACS-X). To fit the plane, we used the Python package LtsFit (ATLAS_XV_Michele_2013), a robust least-squares regression with uncertainties in all dimensions and intrinsic scatter. We did not clip outliers, but included all data points in the fit. Effective radii were fit in kpc, velocity dispersions in units of 250 $\mathrm{km\ s^{-1}}$, and surface-mass densities in $\mathrm{10^{10}\ M_{\odot}\ kpc^{-2}}$. The samples followed the same MFP with coefficients consistent with those found by Auger_2010. The slopes are close but not identical to those in Eq. ( 17), indicating that the assumption of a constant virial coefficient is not exactly correct, but it is likely to have some dependence on mass or velocity dispersion because elliptical galaxies are not completely scale invariant. We overplot the TDCOSMO sample for illustration and show that it lies along the plane across the sample space occupied by the SL2S and SLACS samples. The fit indicates no intrinsic scatter to the combined sample. As a test, we fitted all combinations of subsamples, and the coefficients are the same within the errors.
@@
Figure 11: Lensing-mass fundamental plane with isothermal mass profile assumption, $\mathrm{\log R_{eff}=a\log\sigma_{1/2}+b\log\Sigma_{1/2,iso}+c}$. The plane is fit to the combined SL2S, SLACS/KCWI, and SLACS-X samples. The TDCOSMO sample is overplotted. Here, $\mathrm{\Sigma_{1/2,iso}=M_{R_{eff}/2,iso}/\pi(R_{eff}/2)^{2}}$ is the average lensing surface-mass density within half the effective radius, assuming the mass density profile is isothermal (i.e., $\gamma_{lens}=2$). $\sigma_{1/2}$ is given in units of 250 $\mathrm{km\ s^{-1}}$, $\mathrm{M_{1/2}}$ in units of $10^{10}\ \mathrm{M_{\odot}}$, and $\mathrm{R_{eff}}$ in kpc. The dotted line shows the one-to-one line.
@@
We also studied for the first time the MFP for a generic power-law profile using the measurements of $\gamma=\gamma_{\rm lens}$ that are now available. The uncertainty on $\Sigma_{1/2}$ is given by
@@
$\delta\log\Sigma_{1/2}=\sqrt{(1-\gamma_{\rm lens})^{2}\left(\frac{\delta R_{\rm eff}}{R_{\rm eff}\ln 10}\right)^{2}+\delta\gamma^{2}\left(\log\frac{R_{\rm eff}}{2R_{E}}\right)^{2}}.$ (18)
[…]
@@
The two planes are both well defined, and there are no significant differences between the subsamples. We quantified any residual differences between the subsamples in the following way. Since the subsamples, except SLACS-X, are too small to robustly determine the slopes, we kept the slopes $a$ and $b$ fixed to the values obtained for the combined sample, and we then fit for the intercept $c$ of each subsample. In the isothermal case, the intercept is $0.451\pm 0.019$, $0.463\pm 0.027$, $0.456\pm 0.027$, and $0.445\pm 0.015$ for the combined, SL2S, SLACS-KCWI, and SLACS-X samples, respectively. They are consistent within the errors. The error on the intercept is largest for the SL2S sample and corresponds to $\pm$5.8% in Reff and $\pm 4.2\%$ in $\sigma_{1/2}$. For the nonisothermal case, the intercept is $0.560\pm 0.020$, $0.579\pm 0.029$, $0.561\pm 0.021$, and $0.548\pm 0.022$ for the samples in the same order as above. Again, they are consistent within the errors. The error is largest for the SL2S sample and corresponds to $\pm$6.9% in Reff and $\pm 5.3\%$ in $\sigma_{1/2}$.
@@
Figure 12: Lensing-mass fundamental plane with a flexible power-law mass profile $\mathrm{\log R_{eff}=a\log\sigma_{1/2}+b\log\Sigma_{1/2}+c}$. The plane is fit to the combined SL2S, SLACS/KCWI, and SLACS-X samples. The TDCOSMO sample is overplotted. Here $\mathrm{\Sigma_{1/2}=M_{R_{eff}/2}/\pi(R_{eff}/2)^{2}}$ is the average lensing surface-mass density within half the effective radius. We updated the SLACS-X lens models to those from T24 with free power-law slopes to compare with SL2S and SLACS/KCWI lens models. $\sigma_{1/2}$ is given in units of 250 $\mathrm{km\ s^{-1}}$, $\mathrm{M_{1/2}}$ in units of $10^{10}\ \mathrm{M_{\odot}}$, and $\mathrm{R_{eff}}$ in kpc. The dotted line shows the one-to-one line.
@@
In conclusion, the MFP analysis shows no evidence of structural differences between the SL2S, SLACS, and TDCOSMO samples. This does not mean that there is no evolution in the characteristics of the galaxies. For example, in Figs. 11 and 12, the SL2S galaxies tend to have smaller effective radii, which is generally consistent with the finding that early-type galaxies have smaller sizes at high redshift than in the local Universe for a given stellar mass (e.g., Vd10). Sonnenfeld15 found that the fraction of dark matter within the inner 5 kpc increases with redshift, and William_2025_dinos_2 also found mild evidence for evolution. We showed, however, that galaxies evolve within the lensing-mass fundamental plane. An example of a mechanism that moves galaxies along the lensing-mass plane is the so-called envelope accretion that was recently considered by Nipoti2025 in addition to the standard major and minor mergers. Envelope accretion describes the case in which diffuse satellites are completely disrupted in the galaxy outskirts. As shown in Fig. 13, the model proposed by Nipoti2025 tantalizingly predicts that envelope accretion causes galaxies to evolve almost exactly along the lensing-mass plane. Similar results were obtained for the generalized mass plane with free mass-density slopes. For comparison, we show in the same figure the predicted evolution of galaxies that grow in mass by a factor of two via either minor merging with a mass ratio 1/10 or by equal-mass major dry merging, as predicted by standard toy models (e.g., Nipoti2025). The galaxy follows paths that lie almost parallel to the isothermal MP in these cases as well. However, they form slightly larger angles with the isothermal MP than the angle that formed by the path predicted by the envelope-accretion model. A full comparison with galaxy evolution models is beyond the scope of this paper and is left for future work.
Figure 13: Evolution in the same parameter space as in Fig. 11 of a reference galaxy that grows by a factor of two in stellar mass according to three different processes: Envelope accretion (green arrow), and major (purple arrow) and minor (gray arrow) dry merging. In all cases, the initial galaxy (red symbol) is assumed to lie on the isothermal MP (dashed line). The final position (blue symbol) of the galaxy in the diagram was calculated using the model described in section 6.2 of Nipoti2025 for envelope accretion and using the equations of section 3.1 of Nipoti2025 for minor (mass ratio $\xi=0.1$) and major (mass ratio $\xi=1$) dry merging, assuming a size-mass relation slope $\beta_{R}=0.76$ (see Nipoti2025).